If the demand function is : 𝑄𝑑 = 1300 − 10𝑃 − 𝑃2 Where P = 20, find the price elasticity of demand (show all working used)  


1
Expert's answer
2022-02-17T04:37:56-0500

Price Elasticity of Demand

"Q_d=1300-10P-P^2"

"P=20"

"Q_d=1300-10(20)-20^2"

"Q_d=700"

"Price\ Elasticity = \frac{\%change\ in\ Quantity}{\%change\ in\ Price}"

To get percentage changes, we assume assume the price changes by an increase and decrease of 20%.

Increase;

new price "=1.2\times 20=24"

"Q=1300-10(24)-24^2"

"Q=484"

"Price\ Elasticity=\frac{(484-700)/700}{(24-20)-20}"

"=\bold{-1.5}"

Decrease;

new price "=0.8\times20=16"

"Q=1300-10(16)-16^2"

"Q=884"

"Price\ Elasticity=\frac{(884-700)/700}{(16-20)/20}"

"=Price\ Elasticity=\frac{(484-700)/700}{(24-20)-20}"

"=\bold{-1.3}"

Price Elasticity of Demand at both instances result to a negative number. The product therefore has an inelastic demand function.